Parth (parthkohli):

OpenStudy user s73 has 11,000 medals. OpenStudy user P has 3,000 medals. Openstudy user s73 gains 25 per day. OpenStudy user P gains 50 per day. Assume that P will reach SmartScore 100 if P equals s73, how much time will it take for P to reach a smartscore of 100?

OpenStudy (lgbasallote):

lol

OpenStudy (inkyvoyd):

But but but medals are not worth the same!

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3,000 +50x = 11,000 + 25x }\)

Parth (parthkohli):

'ASSUME'

OpenStudy (lgbasallote):

Assume that P will reach SmartScore 100 if P equals s73?

OpenStudy (lgbasallote):

^what>

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3,000 + 25x = 11,000 }\) \( \color{Black}{\Rightarrow 25x = 8000}\) \( \color{Black}{\Rightarrow x = 320}\) 320 days. Am I correct?

Parth (parthkohli):

lol P is one year away.

OpenStudy (inkyvoyd):

Wrong, cause medal change rate is not constant.

OpenStudy (lgbasallote):

^exactly

OpenStudy (inkyvoyd):

*medal gain rate

Parth (parthkohli):

inky, it's supposed to be a simple question, not to be complicated by that.

OpenStudy (lgbasallote):

and who the heck can equal s73

OpenStudy (inkyvoyd):

and medals is not worth as much as other medals

Parth (parthkohli):

I can equal s73, so can a64 B)

OpenStudy (inkyvoyd):

We must look at the praticality of a model before the solution of one.

OpenStudy (anonymous):

Over 9000 days.....

OpenStudy (lgbasallote):

s.k didnt even make it

Parth (parthkohli):

1 year more to wait xP

OpenStudy (lgbasallote):

s73, a64 and s.k are the top users :p beat em first

Parth (parthkohli):

I have beaten L.

OpenStudy (lgbasallote):

we're all gonna be 100 by then lol

OpenStudy (lgbasallote):

L is gonna come back from behind someday ;)

OpenStudy (lgbasallote):

when you least expect it

Parth (parthkohli):

ORLY?

OpenStudy (lgbasallote):

just wait till monday -_-

OpenStudy (anonymous):

Hi. You're a genius.

OpenStudy (anonymous):

hahaha woww